# Noemi Walczak Université Paris 1 Panthéon-Sorbonne, Paris, France Kenan Huremović Armando Rungi

### v3

@@
Evaluating the EU Carbon Border Adjustment Mechanism with a Quantitative Trade Modelthanks: We are indebted to Lionel Fontagné for initial discussion. We are also grateful to Mathieu Parenti and Maurizio Zanardi for useful comments.
@@
We analyze the economic and environmental impacts of the European Carbon Border Adjustment Mechanism (CBAM) using a multi-country, multi-sector general equilibrium model with input-output linkages. We quantify the general equilibrium responses of trade flows, welfare, and emissions. To our knowledge, we developed the first quantitative trade model that jointly endogenizes both the Emissions Trading System (ETS) allowance and the CBAM prices. We find that CBAM marginally increases EU Gross National Expenditure (GNE) and shifts trade toward more domestic and cleaner production. Emissions embodied in direct EU imports fall by 4.8%, and by 3% when including indirect emissions —underscoring the importance of accounting for production networks in evaluating policy outcomes.
@@
Keywords: Carbon border adjustment, general equilibrium, trade policy, production networks, international trade, climate change
@@ 1. INTRODUCTION
Despite decades of international climate efforts, greenhouse gas (GHG) emissions have continued to rise throughout the 21st century [14]. Since 2005, a cornerstone of the European Union’s climate policy has been the Emission Trading System (ETS), a market-based mechanism that has proven effective in reducing emissions. While the ETS covers the EU as well as Norway, Iceland, and Liechtenstein through the European Economic Area agreement, its scope remains geographically limited. This creates a risk of carbon leakage — where emissions-intensive production is relocated to countries with laxer climate regulations. To counter this, the EU introduced the Carbon Border Adjustment Mechanism (CBAM), which seeks to equalize the carbon cost between imported and domestically produced goods. Since 2023, the CBAM has been in a transitional phase, requiring firms in selected sectors to report the embodied emissions of their imports without financial obligations. Starting in 2026, importers will be required to purchase CBAM certificates corresponding to the carbon content of covered goods, which should reduce incentives to shift emissions abroad and strengthen the environmental integrity of the EU’s climate strategy. Initially targeting six carbon-intensive sectors, the mechanism is expected to expand to all sectors currently covered by the ETS by 2030. In this paper, we refer to the initial stage as the current CBAM, and to the scenario where the CBAM is applied across all ETS sectors as the full CBAM. As a cross-border policy instrument, the CBAM effectively introduces a wedge between the cost of foreign producers and the price faced by importers, with potential implications for trade flows, input sourcing, and broader macroeconomic outcomes that remain uncertain and difficult to assess ex-ante.
@@
In this paper, we propose a quantitative trade model to evaluate the effects of the CBAM on trade, emissions, and welfare. Our model is rich enough to account for the complexity of the world supply network and heterogeneity in the climate policy across countries. In particular, we incorporate input-output linkages at the country-sector level and endogenously determine national carbon prices [^1] We endogenize the emission prices in countries that have a functioning emission market., thereby capturing all types of instruments that put a price on emissions. We calibrate the model using publicly available data collected from standard sources. We find that the CBAM leads to a decrease in the share of carbon-intensive (dirty) imported intermediate inputs by 1.06% and an increase in non-carbon-intensive (clean) imported inputs and domestic inputs by 0.57% and 0.04%, respectively, suggesting the reallocation in intermediate input flows due to CBAM-induced trade barriers. The CBAM leads to a significant reduction in emissions embodied in direct imports by almost 4.80%. However, this number falls to 3% once we account for the emissions embodied in importers’ supply chains, due to substitution towards non-targeted carbon-intensive inputs along the supply chains. Finally, we find a very small effect on the EU gross national expenditure (GNE), with the introduction of the CBAM leading to an increase of 0.006% in the EU GNE.
@@
To evaluate the effects of the CBAM, we start by building a multi-country, multi-sector general equilibrium model with input-output linkages, building on [7]. In the model, a representative firm from each sector in each country produces output using labor and outputs produced by other sectors as inputs. The production technology exhibits constant returns to scale and, importantly, generates emissions as a by-product. Markets are perfectly competitive, and trade in intermediate inputs at the country-sector level shapes the structure of international supply chains. We model the CBAM as a price levied on the carbon content of selected carbon-intensive imported inputs. Importantly, this price is endogenously determined in the model: it is set to match the domestic price of the carbon content of the same sectoral goods, net of the foreign national environmental taxation. To our knowledge, we are the first to endogenize carbon prices and consequently the CBAM prices within a quantitative trade model.
@@
Intuitively, the introduction of CBAM increases the price of carbon-intensive imports from countries with laxer regulations, prompting a reallocation of demand toward cleaner foreign producers and EU-ETS producers. This shift influences the demand for ETS allowances, thereby affecting their market price and, through the endogenous mechanism, the CBAM price itself. Thereby, in the model, we are capturing not only the direct effects of the policy but also its indirect feedback through emissions markets and allowance pricing.
@@
We calibrate the model using publicly available and easily accessible data. In particular, we source data on bilateral trade flows, gross output, and value added from the OECD Inter-Country Input-Output tables, which collect data from 44 sectors across 32 countries, complemented by a residual Rest of the World. We obtain tariffs from the World Trade Organization (WTO) Integrated Database and Consolidated Tariff Schedule. Finally, we obtain the emissions data from the OECD, complementing it with the data from the European Union Transaction Log (EUTL) database for some countries. In countries with an operational national carbon pricing scheme by 2024, we set the total supply of emissions to match the aggregate quantities we retrieve from the respective databases mentioned above. For countries without a pricing scheme in place, we impose an exogenous carbon price equal to the country-specific Effective Carbon Rate, sourced from the OECD, while the supply of emissions is determined endogenously.
@@
We begin by examining how the CBAM affects international trade patterns, focusing on the following three outcomes: the average share of intermediate inputs the EU imports from abroad, the share of intermediate inputs EU sources domestically, and the size of EU sectors as measured by their average Domar weights in the world economy. Next, we turn to our core findings on the emissions embodied in EU imports and carbon leakage. We conclude by evaluating the impact of the CBAM on gross national expenditure in both the EU and its non-member countries. For each outcome, we compare the effects of the current and full CBAM, and distinguish between two scenarios: (i) a benchmark case in which the CBAM value adjusts with emission prices, referred to as the endogenous CBAM; and (ii) a case where the CBAM value is set according to the current mechanism but remains fixed, referred to as the exogenous CBAM.
@@
We find that the current CBAM leads to roughly a 1.06% decline in the average share of imported carbon-intensive goods directly affected by the measure, encouraging a shift in demand towards cleaner alternatives. Accordingly, importers redirect demand toward non-carbon-intensive foreign input providers as well as domestic inputs. Domestic purchases of both clean and polluting goods grow by over 0.03%, reflecting partial substitution of foreign intermediate goods with locally produced alternatives. The size of dirty EU sectors (measured with Domar weight) decreases on average by 0.03% while the size of clean EU sectors increases by 0.11%. The former reduction is driven by an increase in the price of EU materials, thus, a higher marginal cost for EU producers. Notably, the model captures the general equilibrium impact of the CBAM beyond the sectors it directly regulates. The effects more than double under the full CBAM scenario.
@@
Our key insights relate to the changes in emissions embodied in imports. The analysis distinguishes between direct emissions from imported goods and a broader measure that also includes indirect emissions along the importers’ supply chains. We find that the current CBAM reduces European direct import-related emissions by 4.73% (8.84% under the full CBAM). When indirect imports are considered, the reduction in embodied emissions falls to 2.96% (5.19% under the full CBAM). The gap between the change in emissions embodied in direct and total – direct and indirect – imports reflects the broader substitution effects on the supply chain triggered by the policy. In the first instance, the CBAM directly reduces imports of carbon-intensive goods explicitly targeted by the adjustment mechanism. However, this decline induces substitution towards non-targeted inputs, whose relative price falls and whose demand correspondingly rises. As demand for these substitutes increases, so too does the demand for the inputs required to produce them – including carbon-intensive ones. This reallocation weakens the overall impact of the CBAM by increasing carbon emissions upstream. Although the latter is outside the scope of the present regulation, we argue that it is relevant to understand how to reach policy targets after capturing the full carbon footprints along global supply chains. On a higher level of aggregation, we find that the current CBAM should lead to a reduction of emissions leakage by 0.11%.
@@
Interestingly, at the aggregate level, the current CBAM configuration increases real GNE by 0.006%, reaching 0.04% under the full CBAM scenario. The observed increase in EU welfare is primarily driven by terms-of-trade improvements resulting from a reallocation of demand that favors domestic production. These gains reflect higher export prices relative to imports, allowing the EU to obtain more imports per unit of exports and thus benefit from improved trade conditions. In contrast, extra-EU countries experience a minor decline in GNE, with reductions of 0.008% and 0.02% under the current and full CBAM scenarios respectively.
@@
Intuitively, the effects of CBAM strongly depend on two factors: the emissions intensity of production technologies in source countries and the degree of the EU’s integration into global supply chains. To explore these dimensions, we conduct two types of counterfactual exercises. First, to isolate the role of technology, we vary a measure of production cleanliness in source countries. Second, we vary the degree of the EU’s integration into the global supply network. In both exercises, our primary outcome of interest is import-embodied emissions.
Our results show that cleaner global production technologies reduce the impact of CBAM in a more-than-proportional manner. In contrast, deeper supply chain integration weakens the policy’s effectiveness by limiting the scope for substitution toward domestic inputs. Importantly, these two channels are interrelated: EU firms tend to source more heavily from countries that employ dirtier production technologies.
Related literature
This paper bridges two different fields of literature: the literature on the trade effects of environmental policies, more specifically, the literature on carbon border adjustment, and the literature on production networks.
By extending the input-output model by [7] to incorporate emissions in production, we contribute to the literature that studies the interactions between international trade, the environment, and environmental regulations through general equilibrium trade models ([15, 22, 23, 21, 28]). Yet, differently from our framework, previous models do not fully map the global production networks, either because they assume that producers choose to buy inputs only from the lowest cost supplier – as in the workhorse model by [6] – or because they completely neglect the role of intermediate inputs in production.
There is a strand of the literature on carbon border adjustments largely relying on traditional Computable General Equilibrium (CGE) models [4, 20, 5, 25, 3]. Different from these papers, our quantitative trade model allows us to keep track of the economic mechanisms that generate the main results, avoiding the ”black box” associated with CGEs.
The most related to our paper is the growing literature analyzing CBAM with structural trade models ([29, 8, 19, 12, 18]). The papers closest to ours are: [19, 12]. [19] analyze the effect of the European CBAM on trade and emissions within a general equilibrium model. Their incorporation of country-sectoral production linkages à la [6] allows them to obtain disaggregated changes at the sectoral and country level. They find a negative, despite small, change in global emissions driven by the reallocation of production from carbon-intensive to less polluting countries. However, because their framework follows [6] in assuming that each sector imports a given input only from its lowest-cost supplier, it cuts many potential sourcing links and therefore restricts the extent of reallocation across alternative input sources. By contrast, by mapping the full global input-output network, we are able to quantify the propagation of CBAM through multiple direct and indirect channels. Our findings also complement those obtained by [12]. They combine a structural trade model with product-level and balance sheet microdata, running quantitative simulations comparing carbon leakage and welfare effects of ETS-only versus ETS+CBAM scenarios. Their analysis focuses on French firms and captures firm-level sourcing decisions at both the intensive and extensive margin within a framework of monopolistic competition. They find that the CBAM reduces embedded emissions in French imports by inducing a reallocation of purchases toward cleaner inputs. By contrast, our model abstracts from firm heterogeneity and adopts a higher level of aggregation by focusing on sectors, rather than on firms. At the same time, we extend their analysis in several dimensions. First, we adopt a global country-sector perspective, which allows us to quantify the general equilibrium effects of CBAM across multiple sectors and countries. Second, we move beyond the Armington minimum-cost-supplier structure used in their framework and map the full international input-output network. This allows us to trace the propagation of the CBAM through the whole supply chain and to explicitly account for both direct and indirect (supply-chain-mediated) emissions by capturing the complete carbon footprint of global production. Finally, while all these papers treat the CBAM as an exogenous policy shock, we take into consideration its endogenous nature and its interaction with domestic climate policies by explicitly endogenizing both carbon prices and the CBAM prices.
Our work is also related to the literature on production networks [1, 2, 9]. While our model is fairly standard and aligned with this literature, we are the first to examine the network effects of the carbon border adjustment policy.
Overall, we argue that our framework provides a comprehensive assessment of the economic and environmental implications of CBAM, contributing to a broader discussion on climate policy design in an interconnected global economy.
@@
The remainder of the article is structured as follows. In Section 2 we provide more details about the CBAM. In Section 3, we introduce the theoretical framework, while Section 4 presents the results of the simulation of the policy. Finally, Section 6 concludes and offers some final thoughts on the future of CBAM.

### v4

@@
The Network Effects of the EU Carbon Border Adjustment Mechanism with a Quantitative Trade Modelthanks: We are indebted to Lionel Fontagné for initial discussion. We are also grateful to Mathieu Parenti, Maurizio Zanardi and Alessia Campolmi for useful comments.
@@
We investigate the economic and environmental impacts of the European Carbon Border Adjustment Mechanism (CBAM) using a multi-country, multi-sector general equilibrium model with input-output linkages adapted from [caliendo2022distortions]. We quantify the general equilibrium responses of trade flows, expenditures, and emissions. To our knowledge, we are the first to endogenize both carbon prices and the CBAM price. We find that, when CBAM is at full force, it could contribute to reducing carbon emissions by $5.19\%$, a rate that could have been higher ($8.84\%$) without the moderating role of global production networks. As a consequence of CBAM, we observe substitution effects towards non-targeted carbon-intensive inputs located upstream in the supply network. Notably, CBAM marginally increases EU Gross National Expenditure (GNE) due to changes in the terms of trade, and it shifts sourcing towards domestic and environmentally cleaner inputs. In contrast, extra-EU countries experience only a slight decline in GNE ($0.02\%$) and emissions ($0.11\%$). Finally, we propose counterfactual exercises that show the relative importance of technological progress on economic integration in global supply networks. Nonetheless, we argue that supply-chain-wise policies are useful for capturing the full carbon footprint in production.
@@
Keywords: Carbon border adjustment, general equilibrium, trade policy, production networks, climate change, European integration
@@ 1. Introduction
Despite decades of warnings on climate change and the need for global policies, greenhouse gas (GHG) emissions have continued to rise throughout the 21st century [ipccreport]. Since 2005, a cornerstone of the European Union’s climate policy has been the Emission Trading System (ETS), also known as cap-and-trade, a market-based mechanism that has proven effective in reducing emissions. While the ETS covers the EU as well as other countries in the European Economic Area agreement, its scope remains geographically limited. This creates a risk of carbon leakage, where emission-intensive production is relocated to countries with laxer climate regulations. To counter this, the EU introduced the Carbon Border Adjustment Mechanism (CBAM), which seeks to equalize the carbon cost between imported and domestically produced goods after the introduction of certificates whose price is linked to the weekly average price of the ETS.
Since 2023, the CBAM has been in a transitional phase, requiring firms in selected polluting sectors to report the embodied emissions of their imports without financial obligations. Optimistically starting in 2026, importers would be required to purchase CBAM certificates corresponding to the embedded carbon content of covered goods. The scope is to reduce incentives to shift emissions abroad and strengthen the environmental integrity of the EU climate strategy [^1] One of the criticisms made against the CBAM is that it would require a large information system to collect data about trade partners’ emissions at a level of detail that makes it difficult to implement in a short time. See, for example, [campolmi2024designing]..
Initially targeting six carbon-intensive goods, the mechanism is expected to expand to all sectors currently covered by the ETS by 2030. In this paper, we refer to the initial stage as the reduced CBAM, and to the scenario where the CBAM is applied across all ETS sectors as the full CBAM. As a cross-border policy instrument, the CBAM introduces a wedge between the cost of foreign producers and the price faced by importers, with potential implications for input sourcing, gains from trade, and broader macroeconomic outcomes.
@@
Against this previous background, we adopt a quantitative trade model that is rich enough to account for both the complexity of the global supply network and the heterogeneity in climate policies across trade partner countries. In particular, we incorporate input-output linkages for each country-sector and endogenously determine the carbon prices in countries with a functioning emissions market. We calibrate the model using publicly available data collected from standard sources. We find that when the CBAM is fully implemented, it can lead to a significant reduction in emissions embodied in direct imports by 8.84%. However, when we account for the emissions embodied in importers’ supply chains, this number falls to 5.19% due to substitution towards non-targeted inputs along the supply chains. When we focus on imported intermediates, we find that the CBAM can bring a decrease in the share of carbon-intensive (dirty) imported inputs by about 2.14%, which is compensated by an increase in non-carbon-intensive (clean) imported and domestic inputs by 1.27%, evidently suggesting a reallocation effect through changes in the map of relative costs of global inputs. Notably, similar reallocation effects can be found in the Gross National Expenditures of the new equilibrium, as we detect a small albeit positive increase (0.04%). Ultimately, our findings appear to challenge the notion of a trade-off between eliminating carbon leakage, on the one hand, and gains from welfare and trade, on the other.
@@
We start by building a multi-country, multi-sector general equilibrium model with input-output linkages, building on [caliendo2022distortions]. In the model, a representative firm from each sector in each country produces output using labor and the outputs of other sectors as inputs. Sectors can be located in the home country or in other countries. The production technology exhibits constant returns to scale and, importantly, generates emissions as a by-product. Markets are perfectly competitive, and trade in intermediate inputs at the country-sector level shapes the structure of global supply chains. In this context, we model the CBAM as a price levied on the carbon content of selected imported inputs. Importantly, this price is endogenously determined in the model: it is set to match the domestic price of the carbon content of the same sectoral goods, net of the foreign price abroad. To our knowledge, we are the first to endogenize carbon prices, and consequently the CBAM prices, with a quantitative trade model.
@@
Intuitively, the introduction of CBAM increases the price of carbon-intensive imports from countries with laxer regulations, prompting a reallocation of demand towards cleaner foreign producers and EU-ETS producers. This shift influences the demand for ETS allowances, thereby affecting their market price and, through the endogenous mechanism, the CBAM price itself. Therefore, in the model, we capture not only the direct effects of the policy but also its feedback through emissions markets and allowance pricing. At the same time, our model is designed to capture the indirect effects emerging from the existence of complex production networks. An incentive like CBAM, which works as a wedge on direct imports, also has indirect effects upstream, i.e., when we consider, in turn, the importers’ sourcing combination of clean and dirty inputs.
@@
We calibrate the model using publicly available and easily accessible data. In particular, we source data on bilateral trade flows, gross output, and value added from the OECD Inter-Country Input-Output tables, which collect information on 44 sectors across 32 countries, complemented by a residual category known as the Rest of the World. We obtain tariffs from the World Trade Organization (WTO) Integrated Database and Consolidated Tariff Schedule. Finally, we obtain the emissions data from the OECD Environmental Database, complementing it with the data from the European Union Transaction Log (EUTL) database. In countries with an operational national carbon pricing scheme by 2024, we set the total supply of emissions to match the aggregate quantities we retrieve from the respective databases mentioned above. For countries without a pricing scheme in place, we impose an exogenous carbon price equal to the country-specific Effective Carbon Rate, sourced from the OECD, while the supply of emissions is determined endogenously.
@@
After our exercise, we begin by examining how the CBAM affects international trade patterns, focusing on the following three outcomes: the average share of intermediates the EU imports from abroad, the share of intermediates the EU sources domestically, and the size of EU sectors as measured by their average Domar weights in the world economy. Next, we discuss our core findings on the emissions embodied in EU imports and the issue of carbon leakage. We conclude by evaluating the impact of the CBAM on GNE in both the EU and extra-EU countries. For each outcome, we compare the effects of four different scenarios, derived from a $2\times 2$ set of alternatives intended to provide a clearer picture of the policy effects. First, we examine the case of the reduced CBAM, which focuses on a few polluting sectors, and the full CBAM, which extends the policy to all sectors covered by the ETS. Thus, we distinguish the results of both the reduced CBAM and full CBAM, each with two additional alternative scenarios: (i) a benchmark case in which the CBAM value adjusts with emission prices, referred to as the endogenous CBAM; and (ii) a case where the CBAM value is set according to the same mechanism but remains fixed, referred to as the exogenous CBAM.
@@
Let us consider the endogenous full CBAM as our baseline, because it represents the original policy design by the EU authorities. In this case, we find that the CBAM leads to a roughly 2.14% decline in the average share of imported carbon-intensive goods directly affected by the measure, encouraging a shift in demand toward cleaner alternatives. Accordingly, EU importers redirect demand towards non-carbon-intensive foreign and domestic input providers. Purchases of domestic goods grow by a positive albeit small coefficient (0.10%), reflecting partial substitution between foreign intermediate goods and locally produced alternatives. The size of dirty EU sectors (measured using Domar weights) remains unchanged, while the size of clean EU sectors increases by 0.29%, resulting from higher demand from both the EU and extra-EU countries.
@@
Our key insights relate to the changes in emissions embodied in imports. Let us focus more on this aspect. Our analysis distinguishes between direct emissions from imported goods and a broader measure that also includes indirect emissions embodied in the supply chains of EU importers. We find that the full application of CBAM reduces the emissions of direct imports by 8.84%. When indirect imports are considered, the reduction falls to 5.19%. Initially, the CBAM directly reduces imports of carbon-intensive goods that are explicitly targeted by the adjustment mechanism. However, this decline induces substitution towards non-targeted inputs, whose relative price falls and whose demand rises. As demand for these substitutes increases, so too does the demand for the inputs required to produce them – including carbon-intensive ones. Importantly, reallocation weakens the overall impact of the CBAM because it increases carbon emissions upstream. The latter is an insight for the policymakers, who should consider the whole carbon footprint to evaluate the efficacy of the carbon policy versus its targets.
@@
Interestingly, at the aggregate level, the full CBAM configuration increases real GNE by 0.04%. The observed increase in EU welfare is primarily driven by terms-of-trade improvements resulting from a reallocation of demand that favors domestic production. These gains imply that the EU obtains more imports per unit of exports, thereby benefiting from improved trade conditions. In contrast, extra-EU countries experience a minor decline in real GNE with reductions of about 0.02%.
@@
Intuitively, the effects of CBAM strongly depend on two factors: the emissions intensity of production technologies in source countries and the degree of the EU integration into global supply chains. To better explore these dimensions, we conduct two types of counterfactual exercises. First, to isolate the role of technology, we vary a measure of production cleanliness in source countries. Second, we vary the degree of the EU integration into the global supply network. In both exercises, our primary outcome of interest is import-embodied emissions. Our results show that cleaner global production technologies reduce the impact of CBAM in a manner that is more than proportional. In contrast, deeper supply chain integration weakens the policy’s effectiveness by limiting the scope for substitution towards domestic inputs. Importantly, these two channels are interrelated: EU firms tend to source more heavily from countries that employ dirtier production technologies.
@@
The remainder of the article is structured as follows. In Section 2, we relate to previous research. In Section 3, we provide more details about the CBAM. In Section 4, we introduce the theoretical framework, while Section 5 shows how to calibrate and solve the model. Finally, Section 6 presents the results after the policy simulation and Section 7 concludes and offers final thoughts on the future of CBAM.
2. Related literature
This paper bridges two different fields of literature: the one on the trade effects of environmental policies, more specifically, the literature on carbon border adjustment, and the literature on production networks.
By extending the input-output model by [caliendo2022distortions] to incorporate emissions in production, we contribute to studying the interactions between international trade, carbon emissions, and environmental regulations thanks to a structural general equilibrium model. Previous studies include [duan2021environmental, larch2017carbon, larch2024consequences, korpar2023european, shi2025carbon]. Yet, unlike our framework, previous models do not map the global production network, either because they assume that producers choose to buy inputs only from the lowest-cost supplier – as in the workhorse model by [caliendo2015estimates] – or because they completely neglect the role of intermediate inputs in production.
In this respect, there is a strand of research on carbon border adjustments largely relying on traditional Computable General Equilibrium (CGE) models [bohringer2012alternative, ghosh2012border, bohringer2017targeted, morsdorf2022simple, bellora2023eu]. Unlike previous papers, our quantitative trade model tracks the economic mechanisms that generate the main results, thereby avoiding the rigidity and data intensity associated with CGEs.
The most relevant to our paper is the growing literature analyzing CBAM with structural trade models [sogalla2023unilateral, campolmi2024designing, florez2024eu, coster2024firms, farrokhi2025can]. In particular, the papers closest to ours are [florez2024eu] and [coster2024firms]. [florez2024eu] investigate the effect of the EU CBAM on trade and emissions with a general equilibrium model, while incorporating country-sectoral production linkages à la [caliendo2015estimates]. They find a negative albeit small change in global emissions driven by the reallocation of production from carbon-intensive to less polluting countries. However, because their framework assumes that each sector imports a given input only from its lowest-cost supplier, it restricts the extent of reallocation across alternative input sources. By contrast, by mapping the full global input-output network, we quantify the propagation of CBAM through multiple direct and indirect channels, and we provide some insights on the country-sector-level heterogeneity. Our findings also complement those obtained by [coster2024firms]. They combine a structural model with product-level and balance sheet microdata, running quantitative simulations to compare carbon leakage and welfare effects of ETS-only versus ETS-plus-CBAM scenarios. Their analysis focuses on French firms and captures firm-level sourcing decisions at both the intensive and extensive margin with monopolistic competition. They find that the CBAM reduces embedded emissions in French imports by inducing a reallocation of purchases towards cleaner inputs. By contrast, we abstract from firm heterogeneity and adopt a higher level of aggregation to study a global equilibrium after focusing on substitution effects along supply chains in the presence of heterogeneous sourcing strategies. Notably, while all previous papers treat the CBAM as an exogenous policy shock, we consider its endogenous nature as its specific design is such that it should interact with both EU and extra-EU climate policies, and favor the adoption of less carbon-intensive technologies.
Last but not least, our work also relates to literature on production networks [baqaee2020productivity, baqaee2024networks, carvalho2021supply]. In fact, while our model is fairly standard, we are the first to examine the network effects of a policy, the CBAM, which is crucial for solving international coordination problems in the presence of negative externalities, such as those induced by the use of polluting inputs.
Overall, we argue that our framework provides a comprehensive assessment of the economic and environmental implications of CBAM, but also contributes to a broader discussion on the optimal design of policy tools in an interconnected global economy.